How much energy is released per gram of hydrogen for the following reaction? 

2H2(g) +O2 (g) → 2H2O(g) 


a. -483.6 kJ b. -241.8kJ C. -120.9 kJ d. 60.5 kJ e. 241.8 kJ
Answer #1

6. 2H2(g) + O2(g) ---> 2H2O(g)

DH0rxn = (2*-241.82)-(2*0+1*0)

        = -483.64 kj

2 mol H2 = 2*2 g water = -483.64 kj

energy released per gram hydrogen = -483.64/4 = -120.91 kj

answer: c. -120.91 kj